An entrepreneur starts a small firm with one machine, and a small factory space and employs one person initially. With the aid of appropriate diagrams/table and theory, explain the process of production as he increases his output. Indicate clearly what law governs this process of production, the variable and the fixed factors used by this entrepreneur.
1
Expert's answer
2020-06-09T16:30:01-0400

The theory of production tries to explain the principles by which a business firm decides on how much of each commodity that it sells (outputs), it will produce, and how much of each kind of labour, raw material, fixed capital good, etc., that it employs (inputs) it will use.


The theory involves some of the most fundamental principles of economics.


These principles include the relationship between the prices of commodities and the prices of the productive factors used to produce them and also the relationships between the prices of commodities and productive factors, as well as the quantities of these commodities and productive factors that are produced or used.


An example of a production function might be( Q=K+L)


where (Q )is the quantity of output,


( K ) is the amount of capital,


(L) is the amount of labor used in production.


NOTE: This production function says that a firm can produce one unit of output for every unit of capital or labor it employs.

The production function is expressed in the by the formula Q = f(K, L, P, H),


whereby, the quantity produced is a  function of the combined input amounts of each factor.


The fomula for this form is: Q = f(L, K),


In which labor and capital are the two factors of production  with the greatest impact on the quantity of output.